# Deterministic Legal Retrieval: An Action API for Querying the SAT-Graph RAG

### v1

@@ 3.3 API Action Specification
A key design challenge is to provide a search interface that is both simple enough to replicate the behavior of a standard, "temporally-naïve" RAG system, while also exposing the full power of the underlying temporal graph. Traditional RAG pipelines typically index a single, current version of a text. A query like “Which social rights are guaranteed?” executes a semantic search over these current texts.
The SAT-Graph, however, contains all historical versions. A simple semantic search would return a mix of current and outdated, yet relevant, texts. To manage this complexity, the API’s primary search primitive, searchTextUnits, is explicitly designed to handle two core user intents through a simple timestamp parameter:
@@
This design provides a single, powerful action that elegantly bridges the gap between simple RAG queries and complex historical analysis.
3.3.1 Discovery and Search Actions
These actions serve as the primary entry points for an agent, used when it needs to find entities based on ambiguous natural language references or semantic content.
@@ 3.3.1 Discovery and Search Actions
      • reference_text: A string such as "Article 5 of the Constitution", "the Civil Code", or a relative reference like "the previous section".
@@
    • Discussion: This action is indispensable for grounding agentic workflows. Document/component references are often highly ambiguous ("Article 1" could exist in hundreds of laws) or relative ("the following paragraph"). This function handles this by returning a ranked list of candidates rather than forcing a single, potentially incorrect, guess. The optional context_id parameter significantly enhances its precision, enabling sophisticated, context-aware interactions. This design makes the entire system safer and more robust, preventing critical errors that arise from incorrect initial grounding.
@@
    • Discussion: This action is the conceptual counterpart to resolveItemReference. It performs entity linking for thematic nodes, allowing an agent to deterministically identify a theme mentioned by name before using it as a scope in other actions like searchTextUnits.
@@
      • item_ids: (Optional) A list of Item ids. Defines the structural scope. The search includes these items and their descendants.
      • theme_ids: (Optional) A list of Theme ids. Expands the search scope to include all member Items of the specified themes and all of their descendant themes. For example, searching within the "Tribute Law" theme will automatically include items from its sub-themes like "Federal Taxes" and "State Taxes".
@@
      • timestamp: (Optional) A date to constrain the search when using item_ids or theme _ids. If omitted, defaults to the current date ("now"). This parameter is ignored if version_ids is provided.
@@
    • Discussion: This action’s dual-mode design for scoping provides maximum flexibility.
      When to use Mode 1 (Explicit Version Scope): Use when you have already identified a specific set of Versions (e.g., from ‘getItemHistory‘) and need to search exclusively within that curated set. This provides a direct, high-performance path for advanced analytical tasks.
      When to use Mode 2 (Conceptual Scope): Use for natural language queries like "What did Law X say about topic Y in 2005?" The agent defines the "what" (item_ids/theme_ids) and the "when" (timestamp), and the API handles temporal resolution automatically.fic set of Versions of interest (e.g., from getItem History) and needs to perform a semantic search exclusively within that curated set. Using this mode provides a direct, high-performance path for advanced analytical tasks.
      Furthermore, the aspects parameter enables a truly holistic search across the entity hierarchy. By default, the search targets the "canonical" text of the relevant Versions. However, by specifying other aspects (e.g., ["canonical", "textual_metadata"]), the search space expands to include all TextUnits matching those aspects, regardless of whether they are attached to the filtered Versions themselves or to their timeless parent Items. This allows an agent to, in a single query, find a match in a Work’s "publication info." just as easily as in an Version’s official text, creating a comprehensive and context-aware retrieval process.
@@
    • Description: Performs a hybrid search to discover Item nodes (the structural entities of the graph) by combining semantic, lexical, and structured criteria. This action searches across the entire version history of each item’s text to ensure comprehensive discovery.
@@
      • item_ids: (Optional) A list of Item ids to define the search scope.
      • theme_ids: (Optional) A list of Theme ids to expand the search scope.
@@
    • Discussion: This action’s purpose is structural discovery, not content retrieval. When an agent uses searchItems, it is attempting to find the locus of a concept within the graph’s formal structure. To support this, the search must be comprehensive, and is therefore intentionally designed to be time-agnostic, considering content from an item’s entire version history.
@@
      Adding a timestamp to this function would create a confusing overlap and undermine its core purpose as a high-recall structural discovery tool. The item_metadata_filter parameter, in contrast, enhances this purpose by enabling high-precision discovery based on the atemporal metadata of the structural nodes themselves.
@@
    searchThemes(semantic_query?: string, lexical_query?: string,
    top_k?: int) $\rightarrow$ list[{theme: Theme, score: float}]
    • Description: Performs a hybrid search to discover Theme nodes based on a descriptive query. This action searches exclusively on the textual descriptions associated with the themes themselves, not on the content of their member items.
@@
      • top_k: (Optional) The maximum number of candidate themes to return.
    • Returns: A ranked list of candidate objects, each containing a theme’s id and a confidence score.
    • Discussion: This action’s purpose is to map a user’s description of a topic to the most appropriate formal concept in the graph’s ontology. To ensure a precise and high-performance conceptual match, the search is intentionally scoped to the curated textual descriptions of the Theme nodes. Searching the content of all member Items would introduce significant semantic noise and lead to inaccurate results, as large themes could overshadow more specific ones. This design choice maintains a clear separation of concerns: use searchThemes to find the right conceptual category, then use searchTextUnits with the resulting theme’s id to find relevant content within that category.
3.3.2 Deterministic Fetch Actions
Once an agent has acquired a specific ID, these actions are used to retrieve the full, corresponding data objects in a deterministic, non-ambiguous way.
@@ 3.3.2 Deterministic Fetch Actions
    getItem(id: ID) $\rightarrow$ Item
    • Description: Retrieves the full Item object given its unique id.
    • Parameters:
      • id: The canonical id of the item.
  2. 7.
    getTheme(id: ID) $\rightarrow$ Theme
    • Description: Retrieves the full Theme object given its unique id.
    • Parameters:
      • id: The canonical id of the theme.
  3. 8.
    getAction(id: ID) $\rightarrow$ Action
    • Description: Retrieves the full Action data object given its unique identifier. This provides access to all details of a action, including its precise ‘source_version_id‘.
    • Parameters:
      • id: The unique identifier of the action.
    • Returns: The corresponding Action data object.
  4. 9.
    getValidVersion(item_id: ID, timestamp: date-time, policy:
    TemporalPolicy) $\rightarrow$ Version
    • Description: The core temporal resolution function. It finds the single Version that was legally in force on a specific date.
@@
      • policy: An enum specifying the resolution logic, e.g., ‘TemporalPolicy.SnapshotLast‘ which finds the version whose validity interval contains the timestamp.
@@
    • Discussion: The explicit ‘policy‘ parameter is essential for guaranteeing determinism. It allows the agent to specify the exact temporal logic required, such as retrieving the last valid version within a day (snapshot semantics) or handling boundary conditions in a predictable manner. The policy parameter controls temporal resolution: PointInTime: Returns the version valid at exactly the specified timestamp. SnapshotLast: Returns the last known version before or at the timestamp. Default: PointInTime.
  5. 10.
@@
    • Discussion: This action is fundamentally different from searchTextUnits. While the latter is a probabilistic discovery tool based on semantic similarity, getTextForVersion is a deterministic fetch operation. It serves the critical use case where an agent has already identified a specific Version (e.g., via getValidVersion or getItemHistory) and requires its exact textual content. Maintaining this as a distinct, highly-optimized action ensures API clarity, guarantees determinism, and provides a clear audit trail of the agent’s execution plan.
3.3.3 Structural Navigation Actions
These actions are used to traverse the atemporal, hierarchical, and thematic structures of the graph.
  1. 11.
    enumerateItems(item_ids?: list[ID], theme_ids?: list[ID],
    depth?: int) $\rightarrow$ list[Item]
    • Description: Enumerates the structural or thematic members of a given scope. This is the primary action for navigating the graph’s hierarchy and membership relationships.
@@ 3.3.3 Structural Navigation Actions
      • item_ids: (Optional) A list of Item ids. Defines the structural scope. The search includes these items and their descendants.
      • theme_ids: (Optional) A list of Theme ids. Expands the scope to include all member items of these themes.
      • depth: (Optional) An integer controlling the traversal depth for hierarchical Item scopes. If ‘depth=1‘, only direct children are returned. If omitted or negative, all descendants are returned recursively. This parameter is ignored for Theme scopes.
    • Returns: A list of Item objects. The behavior is polymorphic based on the type of the input ‘scope_id‘:
      • For a structural Item scope: it returns its hierarchical children/descendants.
      • For a Theme scope: it returns all of its member Items.
    • Discussion: This action’s purpose is structural enumeration and navigation, not search filtering. It is the correct tool for tasks like building a table of contents or programmatically exploring the members of a conceptual theme. For performing content-based searches within a scope, agents should pass the scope’s id directly to the item_ids or theme_ids parameters of the search functions (searchTextUnits, searchItems), as this is significantly more performant than retrieving the members first and then passing a long list of IDs.
  2. 12.
    getAncestors(item_id: ID) $\rightarrow$ list[Item]
    • Description: Navigates the hierarchy upwards from a specific structural Item to its parent, grandparent, and so on, up to the root Work. This is crucial for retrieving the full contextual path of any given provision.
@@
    • Returns: An ordered list of Item objects representing the hierarchical path, from the highest-level ancestor (e.g., a ‘Title‘) down to the immediate parent of the input item. The root Work itself is typically excluded.
    • Discussion: This action is the fundamental tool for contextualization. The legal meaning of a provision is often determined by the Chapter or Title it belongs to. After a search action like searchTextUnits identifies a relevant item, an agent should use getAncestors to retrieve its structural "breadcrumb". This allows the generation of richer, context-aware responses, significantly improving the quality and explainability of the final answer.
  3. 13.
@@
    • Discussion: This action is the essential counterpart to using a Theme as a scope in enumerate Items. It makes the relationship between items and themes fully bidirectional, enabling an agent to efficiently answer questions like "What topics does this article cover?". Without this primitive, an agent would need to resort to a highly inefficient brute-force search across the entire thematic ontology. This function completes the API’s structural navigation toolkit.
  4. 14.
    getItemContext(item_id: ID) $\rightarrow$ StructuralContext
    • Description: Retrieves the complete, immediate structural context of a single Item in one efficient call. This includes its parent, siblings, and direct children.
    • Parameters:
      • item_id: The id of the Item for which to retrieve the context.
    • Returns: A StructuralContext object containing the full data for the target item and its immediate hierarchical neighbors.
      • target (Item): The full object of the item in question.
      • parent (Item)?: The full object of its immediate parent.
      • siblings (list[Item])?: A list of full objects for its siblings (other items with the same parent).
      • children (list[Item])?: A list of full objects for its immediate children.
    • Discussion: This is a high-utility, performance-oriented action designed for applications that need to render a contextual view of an item (e.g., a document browser or a table of contents). Replicating this functionality would require multiple, less efficient API calls to fetch ancestors, descendants, and then compute the sibling set. By providing this as a single, atomic operation, the API dramatically simplifies the logic for common UI and agentic navigation tasks.
3.3.4 Causal & Lineage Analysis Actions
These actions enable agents to traverse the temporal and causal dimensions of the SAT-Graph, understanding the history and evolution of legal items.
  1. 15.
@@ 3.3.4 Causal & Lineage Analysis Actions
    • Description: Provides a complete, ordered timeline of all actions that have affected a given Item.
@@
    • Discussion: This action provides a complete, auditable narrative of an item’s evolution. An agent can iterate through the returned list of Actions to reconstruct every state transition, accessing the "before" state via ‘action.terminates_version‘ and the "after" state via ‘action.produces_version‘ for each step.
  2. 16.
@@
    • Description: Traces the direct causal links for a single Version. It identifies the specific Action that brought this Version into existence (its provenance) and, if applicable, the Action that terminated its validity.
@@
    • Discussion: This action is crucial for pinpointing the exact legislative act responsible for a specific version of a legal text, facilitating precise auditability.
  3. 17.
    getVersionsInInterval(item_id: ID, start_date: date,
@@
    • Description: Retrieves all Versions of a given Item that were legally valid at any point within a specified time interval.
@@
      • item_id: The id of the item of interest.
      • start_date: The start date of the time interval for the query.
@@
    • Discussion: This function serves as a critical performance optimization for historical analysis. While getItemHistory retrieves the causal events (‘Actions‘), this action retrieves the resulting states (‘Versions‘) within a specific temporal window. Replicating this on the client-side would require fetching the entire history and performing complex filtering. By providing this as a server-side primitive, the API enables an agent to efficiently answer questions like "show all versions of this article during a specific presidential term," which is essential for building responsive and scalable legal research applications.
  4. 18.
@@
    • Discussion: This action provides a powerful, server-side analysis that goes beyond a simple textual ‘diff‘. Because the server has access to the full graph structure, it can perform a structural comparison, identifying not just changed words but also added or removed components (e.g., a new paragraph). This transforms the comparison from a mere text-level operation into a true structural impact analysis, which is fundamental for legal practitioners. Centralizing this complex logic ensures high performance and a consistent, high-quality analysis for all client applications.
  5. 19.
@@
    • Description: An essential action for impact analysis that provides forward causal tracing. Given the id of a source Work (e.g., an amending law), it retrieves all atomic Actions authorized by any clause within that Work.
@@
    • Discussion: This action is the logical inverse of traceCausality. While the latter traces backwards from a single effect to its granular cause, getActionsBySource traces forwards from a high-level cause (an entire legislative document) to all of its granular effects. It is the primary, high-performance primitive for answering the critical question: "What did Law X do?". The optional action_types filter allows for more targeted impact analysis, such as isolating only the revocations or creations enacted by a specific law.
3.3.5 Macro-Level Analysis Actions
Beyond entity retrieval and causal tracing, the Canonical Action API provides a distinct class of actions designed for high-level, aggregate analysis. These functions represent a deliberate architectural choice to offload computationally intensive processing from the client to the server. While an agent could theoretically replicate their logic by making numerous, sequential calls to lower-level primitives, doing so would be highly inefficient and impractical for large-scale analysis.
By treating server-side aggregation as a core primitive, the API enables agents and applications to efficiently answer broad, macro-level questions about legislative evolution. Actions in this category typically do not return lists of full entity objects, but rather structured, lightweight summary objects (like an ImpactReport) containing aggregate statistics and lists of identifiers, designed to be selectively hydrated by the agent as needed.
  1. 20.
    summarizeImpact(item_ids?: list[ID], theme_ids?: list[ID],
    time_interval: TimeInterval, action_types?: list[string]) $\rightarrow$
    ImpactReport
    • Description: An aggregate analysis action that provides a high-level summary of changes that occurred within a specific scope during a defined time period.
    • Parameters:
      • item_ids: (Optional) A list of Item ids to define the structural scope of the analysis.
      • theme_ids: (Optional) A list of Theme ids to define the thematic scope of the analysis.
      • time_interval: A start and end date for the analysis window.
      • action_types: (Optional) A list of strings to filter the summary to specific action types (e.g., [‘Revocation‘, ‘Creation‘]). If omitted, all action types are included.
    • Returns: A structured ImpactReport object, containing aggregate statistics (e.g., counts of each action type) and lists of identifiers for the relevant Actions and affected Items.
    • Discussion: This powerful action enables macro-level analysis and is a deliberate design choice to offload complex processing to the server. Replicating its logic on the client-side would require numerous, inefficient calls. By treating server-side aggregation as a core primitive, the API ensures efficiency and simplifies agent logic for this common legislative analysis use case. The ImpactReport intentionally contains lightweight identifiers, allowing the agent to selectively hydrate full details using batch operations like getBatchActions as needed.
3.3.6 Introspection & Metadata Actions
To enable agents to be more adaptive and robust, the API provides a set of introspection actions that allow them to discover the capabilities and boundaries of the knowledge graph at runtime.
  1. 21.

### v2

@@ 3.3 Canonical Primitive API Specification
A key design challenge is to provide a search interface that is both simple enough to replicate the behavior of a standard, "structurally and temporally-naïve" RAG system, while also exposing the full power of the underlying temporal graph. Traditional RAG pipelines typically index a single, current version of a text. A query like “Which social rights are guaranteed?” executes a semantic search over these current texts.
The SAT-Graph, however, contains all historical versions. A simple semantic search would return a mix of current and outdated, yet relevant, texts. To manage this complexity, the API’s primary search primitive, searchTextUnits, is explicitly designed to give the agent precise control over temporal scope through a simple timestamp parameter:
@@
This design provides a single, powerful primitive that empowers an agent to seamlessly switch between simple, RAG-like queries and complex historical analysis, depending on its strategic goal.
3.3.1 Discovery and Search Primitives
These primitives serve as the primary entry points for an agent, used when it needs to find entities based on ambiguous natural language references or semantic content.
@@ 3.3.1 Discovery and Search Primitives
      • reference_text: A string such as "Article 5 of the Constitution" or "Civil Code".
@@
    • Discussion: This primitive is indispensable for grounding agentic workflows. Document/component references are often highly ambiguous ("Article 1" could exist in hundreds of laws). This function handles this by returning a ranked list of candidates rather than forcing a single, potentially incorrect, guess. The optional context_id parameter significantly enhances its precision, enabling sophisticated, context-aware interactions. This design is a cornerstone of agent safety. By forcing the agent to explicitly handle a list of candidates and their confidence scores, it prevents catastrophic errors that arise from incorrect initial grounding. This initial, auditable step of uncertainty management is what makes subsequent deterministic plans trustworthy.
@@
    • Discussion: This primitive is the conceptual counterpart to resolveItemReference. It performs entity linking for thematic nodes, allowing an agent to deterministically identify a theme mentioned by name before using it as a scope in other primitives like searchTextUnits.
@@
      • item_ids: (Optional) A list of Item ids. The search is strictly constrained to the provided items. This primitive does not automatically expand to include descendant items; if hierarchical search is desired, the agent is responsible for first constructing the full list of item IDs using a navigation primitive like getItemHierarchy.
      • theme_ids: (Optional) A list of Theme ids. The search is strictly constrained to Items that are direct members of the provided themes. This primitive does not automatically expand to include descendant themes; if hierarchical search is desired, the agent is responsible for first constructing the full list of theme IDs using a navigation primitive like getThemeHierarchy.
@@
      • timestamp: (Optional) An ISO 8601 date-time string to constrain the search. If omitted, the search defaults to the current UTC timestamp at the moment of query execution, effectively retrieving the "presently valid" texts. This parameter is ignored if version_ids is provided.
@@
    • Discussion: This primitive is unique within the API as it bridges probabilistic discovery and deterministic scoping. While it can accept formal IDs (version_ids), its primary purpose is to perform a search, which is inherently probabilistic. This is an intentional design choice to empower the agent with a powerful tool for focused analysis. This hybrid behavior is critical for sophisticated reasoning. It allows an agent to first build a high-confidence, verifiably correct set of documents (e.g., "all versions of this law during a specific presidential term") and then perform a semantic exploration within that trusted context. This is fundamentally different from a deterministic fetch like getBatchTexts, which would return all texts for the given versions without any semantic ranking or scoring.
@@
    • Description: Performs a hybrid search to discover Item nodes (the structural entities of the graph) by combining semantic, lexical, and structured criteria. This primitive searches across the entire version history of each item’s text to ensure comprehensive discovery.
@@
      • item_ids: (Optional) A list of Item ids. The search is strictly constrained to the provided items. This primitive does not automatically expand to include descendant items; if hierarchical search is desired, the agent is responsible for first constructing the full list of item IDs using a navigation primitive like getItemHierarchy.
      • theme_ids: (Optional) A list of Theme ids. The search is strictly constrained to Items that are direct members of the provided themes. This primitive does not automatically expand to include descendant themes; if hierarchical search is desired, the agent is responsible for first constructing the full list of theme IDs using a navigation primitive like getThemeHierarchy.
@@
    • Discussion: This primitive is a critical tool for an agent’s grounding strategy. Its purpose is structural discovery, not content retrieval. When an agent uses searchItems, it is attempting to find the timeless locus of a concept within the graph’s formal structure. This stable, atemporal anchor is the necessary prerequisite for forming deterministic, time-bound queries later on. To support this, the search must be comprehensive, and is therefore intentionally designed to be time-agnostic, considering content from an item’s entire version history.
@@
3.3.2 Deterministic Fetch Primitives
Once an agent has successfully grounded a query to a specific ID, it transitions from probabilistic discovery to a fully deterministic execution path. The following primitives form the core of this path, enabling the agent to retrieve data objects with absolute, non-ambiguous certainty, forming the verifiable steps in its audit trail.
@@ 3.3.2 Deterministic Fetch Primitives
    getTheme(id: ID) $\rightarrow$ Theme
    • Description: Retrieves the full Theme object given its unique identifier.
@@
      • id: The canonical identifier of the theme.
    • Discussion: This primitive exemplifies the API’s core pattern for deterministic, primary-key-based retrieval. For brevity and clarity, we detail only getTheme, but the API provides analogous fetch primitives for all other core data models, including getItem(id: ID), getVersion(id: ID), and getAction(id: ID). These functions are the fundamental tools an agent uses to "hydrate" a referenced ID into its full, verifiable data object, forming a critical step in any reasoning plan.
@@
    getValidVersion(item_id: ID, timestamp: date-time) $\rightarrow$ Version
    • Description: The core deterministic temporal resolution function. It answers the critical question of ’what was the legislation on this exact date?’ by retrieving the single Version that was legally in force.
@@
  3. 7.
@@
    • Discussion: This primitive is fundamentally different from searchTextUnits. While the latter is a probabilistic discovery tool based on semantic similarity, getTextForVersion is a deterministic fetch operation. It serves the critical use case where an agent, having already identified a specific Version through a deterministic navigation or temporal query, requires its exact textual content. This primitive represents the final step in many reasoning plans: the retrieval of the ground-truth text. Maintaining it as a distinct, primary-key-based primitive is fundamental to the API’s design, ensuring that this final, crucial step is guaranteed to be deterministic and auditable.
3.3.3 Structural Navigation Primitives
These primitives equip an agent with the tools to deterministically navigate the graph’s timeless structures. Unlike search primitives that discover what is relevant, these primitives allow the agent to understand where a piece of information is located and how it relates to its neighbors. This capability is fundamental for constructing context-aware and legally precise reasoning plans.
  1. 8.
    getItemHierarchy(item_id: ID, depth?: int) $\rightarrow$ list[ID]
    • Description: Allows an agent to systematically discover the structural descendants of a known Item. This is the primary deterministic primitive for exploring the graph’s structural hierarchy.
@@ 3.3.3 Structural Navigation Primitives
      • item_id: The ID of the root Item for the hierarchy traversal.
      • depth: (Optional) An integer controlling the traversal depth. If ‘depth=1‘, only direct children are returned. If omitted or negative, all descendants are returned recursively.
    • Returns: A list of descendant Item IDs.
    • Discussion: This primitive’s purpose is purely structural enumeration. It provides the agent with a deterministic tool for exploration, distinct from probabilistic search. For clarity and brevity, this specification details the hierarchy traversal for Item entities. The same pattern is applied consistently across the API for other hierarchical structures, with analogous primitives such as getVersionHierarchy, getThemeHierarchy and getItemTypeHierarchy available for traversing their respective taxonomies.
  2. 9.
    getItemAncestors(item_id: ID) $\rightarrow$ list[Item]
    • Description: Provides an agent with the full structural lineage of an item, navigating upwards to the root Work. This primitive is fundamental for contextual grounding.
@@
    • Returns: An ordered list of Item objects representing the hierarchical path, from the highest-level ancestor (e.g., a ‘Title‘) down to the immediate parent of the input item. The root Work itself is excluded.
    • Discussion: This primitive is the fundamental tool for contextualization. The legal meaning of a provision is often determined by the Chapter or Title it belongs to. After a search primitive identifies a relevant item, an agent must use getItemAncestors to retrieve its structural ’breadcrumb’. This is not just for generating richer responses; it is a core requirement for explainable reasoning, as the context provided by the ancestors is often part of the answer itself.
  3. 10.
@@
    • Discussion: This primitive makes the relationship between items and themes fully bidirectional, enabling an agent to efficiently answer the essential question: ’What are the official conceptual classifications of this law?’ Without this primitive, an agent would need to resort to a highly inefficient brute-force search across the entire thematic ontology. This function completes the API’s structural navigation toolkit.
3.3.4 Causal & Lineage Analysis Primitives
This category of primitives moves beyond static navigation to empower an agent with forensic analysis capabilities. They are the tools for traversing the temporal and causal dimensions of the graph, enabling the agent to deterministically reconstruct the history and evolution of any legal provision.
  1. 11.
@@ 3.3.4 Causal & Lineage Analysis Primitives
    • Description: Provides an agent with the complete, unbroken chain of custody for a legal provision, returning an ordered timeline of every Action that has ever affected it.
@@
    • Discussion: This primitive is the foundation of the agent’s auditability. It provides a complete, verifiable narrative of an item’s evolution. By iterating through the returned list of Action objects, the agent can deterministically reconstruct every past state of the provision, making its historical reasoning transparent and defensible.
  2. 12.
@@
    • Description: Allows an agent to perform pinpoint causal attribution for a single Version. It deterministically identifies the specific Action that created it (its provenance) and, if applicable, the Action that ended its validity.
@@
    • Discussion: For an agent, this is the tool for answering the ’why’ and ’how’ of legal change. It is crucial for pinpointing the exact legislative act responsible for any given state of the law, forming the core of a truly explainable and auditable reasoning process.
  3. 13.
    getVersionsInInterval(item_ids?: list[ID], start_date: date,
@@
    • Description: Retrieves all Versions of the given Items that were legally valid at any point within a specified time interval.
@@
      • item_ids: (Optional) A list of Item ids. The search is strictly constrained to the provided items. This primitive does not automatically expand to include descendant items; if hierarchical search is desired, the agent is responsible for first constructing the full list of item IDs using a navigation primitive like getItemHierarchy.start_date: The start date of the time interval for the query.
@@
    • Discussion: This function serves as a critical performance optimization for historical analysis. While getItemHistory retrieves the causal events (‘Actions‘), this primitive retrieves the resulting states (‘Versions‘) within a specific temporal window. Replicating this on the client-side would require fetching the entire history and performing complex filtering. By providing this as a server-side primitive, the API enables an agent to efficiently answer questions like "show all versions of this article during a specific presidential term," which is essential for building responsive and scalable legal research applications.
  4. 14.
@@
    • Discussion: This primitive provides a powerful, server-side analysis that goes beyond a simple textual ‘diff‘. Because the server has access to the full graph structure, it can perform a structural comparison, identifying not just changed words but also added or removed components (e.g., a new paragraph). This transforms the comparison from a mere text-level operation into a true structural impact analysis. For an agent tasked with explaining a change, this is a critical capability. Centralizing this complex logic on the server provides the agent with a consistent, high-performance tool for generating precise and reliable change summaries.
  5. 15.
@@
    • Description: An essential primitive for impact analysis that provides forward causal tracing. Given the id of a source Work (e.g., an amending law), it retrieves all atomic Actions authorized by any clause within that Work.
@@
    • Discussion: This primitive is the logical inverse of traceCausality. While the latter traces backwards from a single effect to its granular cause, getActionsBySource traces forwards from a high-level cause (an entire legislative document) to all of its granular effects. It is the primary, high-performance primitive that empowers an agent to answer the critical question of impact analysis: ’What were all the concrete effects of Law X?’. The ability to trace causality forward from a single legislative source to all its distributed effects is a sophisticated reasoning capability now made available as a deterministic primitive.
3.3.5 Introspection & Metadata Primitive
For an agent to be truly autonomous and trustworthy, it must be aware of its environment’s capabilities and limitations. This class of primitives provides this essential capability—a form of programmatic self-awareness. They allow an agent to discover the boundaries of the knowledge graph at runtime, enabling it to formulate more robust plans and interact more intelligently with the user.
  1. 16.